# Unsourced Random Access: A Comprehensive Survey

### v2

@@ Abstract
Multiple access communication systems enable numerous users to share common communication resources simultaneously, playing a crucial role in wireless networks. With the emergence of the sixth generation (6G) and beyond communication systems, supporting massive machine-type communications with sporadic activity patterns is expected to become a critical challenge. Unsourced random access (URA) has emerged as a promising paradigm to address this challenge by decoupling user identification from data transmission through the use of a common codebook. This survey provides a comprehensive overview of URA solutions, covering both theoretical foundations and practical implementations. We present a systematic classification of URA solutions across three main channel models: Gaussian multiple access channels (GMACs), single-antenna fading, and multiple-input multiple-output (MIMO) fading channels. For each category, we analyze and compare state-of-the-art solutions in terms of performance, complexity, and practical feasibility. Additionally, we discuss critical challenges such as interference management, computational complexity, and synchronization issues. The survey concludes with promising future research directions and potential methods to address existing limitations, providing a roadmap for researchers and practitioners in this rapidly evolving field.
Index Terms:
Multiple access, RA techniques, grant-free, unsourced RA, machine-type communications, Internet of things
@@ I Introduction
The rapid proliferation of Internet of Things (IoT) devices is driving a fundamental shift in wireless communication systems. Current estimates project over 15 billion connected IoT devices by 2023, with the expected increase to a staggering number of over 38 billion devices by 2029 [1]. One example of such systems is machine-type communications (MTC) (also known as machine-to-machine (M2M) communications), which, as a key aspect of the IoT paradigm, enables various machines, sensors, and devices to exchange information without direct human intervention. MTC plays a crucial role in many applications ranging from industrial automation and smart cities to healthcare monitoring and environmental sensing [2]. Specifically, massive MTC is anticipated to accommodate a density of up to one million devices (per square kilometer), which typically operate with low computational and storage capabilities, low duty cycles, and small payloads [3]. This massive scale presents unprecedented challenges for wireless access, as traditional communication paradigms were designed for a relatively small number of high-data-rate users rather than an enormous number of devices with sporadic, low-rate transmissions.
Multiple access (MA) techniques, which enable different users to share a common communication medium simultaneously, have been fundamental to wireless communication systems [4]. MA technologies can be broadly categorized into centralized access and random access (RA). In centralized access (also known as coordinated access), a central authority controls and manages access to the shared communication channel by allocating distinct resources to each user for data transmission. This approach has been the predominant MA mode in the fifth generation (5G) and previous cellular technologies. However, centralized coordination becomes impractical for massive IoT scenarios due to prohibitive signaling overhead and coordination complexity. In contrast, in RA (also known as uncoordinated access), users transmit their messages simultaneously as soon as new data is generated, without requiring explicit coordination with the central unit. This approach reduces both latency and signaling overhead. These advantages make RA solutions particularly suitable for next-generation (6G and beyond) communication networks, especially when serving a large number of users [5, 6].
@@
Traditional RA solutions like ALOHA [7] and its variants suffer from excessive collisions as the number of users grows. More sophisticated grant-free RA schemes [8] employ unique signatures or preambles for each potential user to facilitate collision reduction at the physical layer. Grant-free RA solutions can accommodate more users compared to ALOHA, as they can employ user scheduling via a limited amount of feedback [9], or device-specific patterns such as codebook structures and spreading sequences [10]. However, the distinct pilot/preamble requirement for each potential user becomes infeasible as the number of connected devices reaches hundreds of millions. On the other hand, in many of the emerging applications, such as such as massive MTC and IoT, devices become active only sporadically. Specifically, massive MTC is anticipated to accommodate a density of up to one million devices (per square kilometer) with low computational and storage capabilities, low duty cycles, and small payloads [3].
@@
To mitigate the challenge of identifying and decoding messages from such a large number of potential devices, Polyanskiy introduced unsourced random access (URA) paradigm [11]. URA fundamentally reimagines the massive access problem by decoupling user identification from data transmission. Instead of assigning unique preambles or codebooks to each potential user employ a common codebook, and the receiver focuses solely on recovering the list of transmitted messages. This paradigm shift drastically reduces system complexity and enables practical implementations for truly massive numbers of devices. Following Polyanskiy’s pioneering work, several low-complexity coding solutions have been developed in the subsequent literature, in which, by employing successive interference cancellation (SIC) techniques, collisions are resolved up to a certain level depending on user transmit powers. Many techniques, such as random spreading [12], multistage orthogonal pilots [13], and time frame slotting [14], have been employed to mitigate the collisions and interference. While a single-antenna receiver in a Gaussian MA channel (GMAC) setup is considered in the earlier works, the literature is shifted to fading channels with multiple-antenna receivers, especially massive multiple-input multiple-output (MIMO), to support more practical channels and an even greater number of users. A block diagram of different MA techniques is depicted in Fig. 1 alongside the three main set-ups in URA research, namely, GMAC, and single-antenna and MIMO fading channels. The figure also shows the way this survey is structured.
Although close to ten years have passed since Polyanskiy’s initial work, and unsourced random access has been studied in many papers, there have been limited efforts to review the existing schemes on URA. The first attempt is a survey on massive access [15], the authors review both many-user MA, where each user is assigned an individual codebook (i.e., the coordinated approach), and URA, which they call uncoordinated massive access. However, since URA was in its early stages at the time of this survey, it mostly covers the earlier works of GMAC URA. In [16], the focus is again on GMAC URA, and the authors review the design paradigms and existing coding schemes for URA over GMAC while fading channels and MIMO URA are presented as open problems. In a less URA-oriented survey [17], an overview of compressive sensing-based grant-free massive access is covered, where some compressed sensing-based URA schemes are also included. In a recent survey [18], through an overview of the existing URA schemes studying mostly GMAC, the authors present URA as a suitable candidate for next-generation communication systems and show that it can support the 3rd Generation Partnership Project (3GPP) standardization.
While the existing surveys focus on the GMAC setup, this survey also covers the fading URA solutions, which comprise the majority of works in the URA literature. More specifically, we conduct a comprehensive review of the existing URA solutions in the literature in its three main setups, namely, URA over GMAC and single-antenna and MIMO fading channels, accompanied by a discussion on the main challenges with possible solutions and many promising future research directions. We study works that put forth coding schemes as well as those that offer fundamental information-theoretical limit calculations for the URA setup. To the best of our knowledge, this is the first comprehensive survey on URA, reviewing this crucial emerging paradigm for 6G and beyond communications in all aspects. A summary of the existing surveys and our contribution can be found in Table I.
@@
IDMA interleave-division MA
@@
MS-MRA multi-stage set-up with multiple receive antennas
@@
MSUG multi-stage set-up with user grouping
@@
The remainder of the survey is organized as follows. In Section II, a general discussion of MA techniques is provided. We start with coordinated MA techniques and find our way to random access. In Section III, we focus on URA as a new paradigm for uncoordinated MA communications with a massive number of potential devices, enumerate its advantages, and hint at some solutions developed in the literature to address the related challenges. Sections IV and V are dedicated to in-depth overviews of the existing URA solutions, with Section IV focusing on URA over GMAC and Section V on URA over fading channels. Regarding communication models over practical wireless channels, we first investigate the fading works with a single antenna receiver and describe the general approaches; then, we discuss those with a MIMO assumption. Finally, we conclude the survey in Section VI with remarks on URA solutions and provide future research directions.
@@ II Background: Multiple Access
Multiple access refers to the scenario in which two or more users simultaneously share the system resources. In this section, we provide a concise overview of the main MA approaches and enumerate their pros and cons in order to better understand the position of URA solutions in next-generation communication systems.
@@ II-A Coordinated Access
One way to serve multiple users is to assign dedicated resources (e.g., time, frequency, code, or their combinations) to them by a central coordination unit through a prior handshaking procedure. The conventional techniques in this framework are time-division MA (TDMA), frequency-division MA (FDMA), orthogonal frequency-division MA (OFDMA), code-division MA (CDMA), and space-division MA (SDMA). In TDMA and FDMA, non-overlapping time slots and frequency sub-channels are assigned to the different users, respectively, and each user utilizes its allocated time frame or frequency band for transmission [4]. On the other hand, in CDMA, all the users can simultaneously utilize all the time-frequency resources through different (nearly) orthogonal code sequences coupled with a low-complexity decorrelation-based receiver. OFDMA provides frequency division multiplexing by assigning different sets of subchannels to the individual users in an orthogonal frequency division multiplexing (OFDM) system, while SDMA relies on the principle of creating (nearly) orthogonal spatial channels for different users to minimize the inter-user interference by employing beamforming and spatial multiplexing [19].
All of the approaches described in the previous paragraph are in the category of OMA, where the number of users is strictly limited due to the orthogonal resource assignment. In order to increase the number of supported users, non-orthogonal MA (NOMA) techniques are investigated in the literature, where multiple users can utilize the same resources to transmit their data, allowing the system to support more users. However, this imposes inter-user interference on the system, and more sophisticated interference-canceling receivers are needed with an increased complexity compared to OMA.
NOMA schemes are divided into two main categories: power-domain NOMA and code-domain NOMA. In power-domain NOMA [20], the users employ different power levels to transmit their data using the same resources, and the receiver exploits the received power differences for detection through successive interference cancellation (SIC). On the other hand, code-domain NOMA is similar to classical CDMA in the sense that the users employ user-specific code sequences to transmit their data. The code sequences are sparse or non-orthogonal sequences with low cross-correlations [21]. Here, the receiver can employ message-passing algorithms utilizing the code structure to differentiate the users. Some examples of code-domain NOMA schemes in the literature are low-density spreading code division MA [22], low-density spreading aided orthogonal frequency division multiplexing [23], sparse code MA (SCMA) [24], and multiuser shared access [25].
@@ II-B Random Access
All of the schemes described in the previous subsection are coordinated. That is, they require scheduling grants and coordination with the BS before transmission, and it is assumed that the number of users and their activity patterns are available at the BS. However, when the number of users becomes large, they become infeasible due to the excessive delays and signaling overhead. For example, the typical number of users per signal dimension in the time-frequency domain (per degree of freedom) is between 1.5-3 according to 3GPP studies, as the system performance significantly degrades beyond that threshold [15]. The common solution addressing these drawbacks is random access, where the users perform their transmission without any coordination by the BS whenever they have data. ALOHA and grant-free random access are the two common RA approaches in the literature.
@@ II-B1 ALOHA & Variants
An early RA approach is the ALOHA protocol [7] (also known as pure ALOHA) proposed in the 1970s, where a device sends the entire packet at a randomly selected time. If no acknowledgment is received within a certain period of time, it is assumed that a collision has occurred with a packet sent by another device, and the packet is retransmitted after an additional random waiting time to avoid repeated collisions. This process is repeated until a successful transmission and its subsequent confirmation occur or until the process is terminated by the user console. Slotted ALOHA (SA) is a variant of ALOHA, where the packets are transmitted in designated time slots, doubling the maximum throughput compared to the pure ALOHA [26]. However, collision is still a major challenge in SA, as several active users may share the same slot to send their data packets, leading to packets colliding within that slot and the discarding of data. The transmission structures in pure ALOHA and slotted ALOHA are illustrated in Fig. 2.
Figure 2: An illustration of pure and slotted ALOHA. Red boxes show the collided packets and green boxes are the successfully transmitted ones.
With the motivation to increase the throughput, several variants of SA are proposed in the recent literature, where the main idea is to transmit multiple copies of a packet in different slots. If it is successfully recovered in one slot, the effect of its replicas in the other slots is removed. Based on this idea, two copies of the same packet is transmitted in contention resolution diversity slotted ALOHA (CRDSA) [27], and the packets are repeated a different number of times rather than twice in irregular repetition slotted ALOHA (IRSA) proposed in [28]. An extension of IRSA exploiting forward error correction that is called coded slotted ALOHA (CSA) [29] approaches a throughput of almost one when sufficiently low coding rates are employed. IRSA and CSA are further studied in [30, 31, 32, 33, 34]. Moreover, the resolution of partial collisions in random access setups is discussed in [35].

### v3

@@ Abstract
Multiple access communication systems enable numerous users to share common communication resources, playing a crucial role in wireless networks. With the emergence of the sixth generation (6G) and beyond communication networks, supporting massive machine-type communications with sporadic activity patterns is expected to become a critical challenge. Unsourced random access (URA) has emerged as a promising paradigm to address this challenge by decoupling user identification from data transmission through the use of a common codebook. This survey offers a comprehensive overview of URA solutions, encompassing both theoretical foundations and practical applications. We present a systematic classification of URA solutions across three primary channel models: Gaussian multiple access channels (GMACs), single-antenna fading channels, and multiple-input multiple-output (MIMO) fading channels. For each category, we analyze and compare state-of-the-art solutions in terms of performance, complexity, and practical feasibility. Additionally, we discuss critical challenges such as interference management, computational complexity, and synchronization. The survey concludes with promising future research directions and potential methods to address existing limitations, providing a roadmap for researchers and practitioners in this rapidly evolving field.
@@ I Introduction
The rapid proliferation of Internet of Things (IoT) devices is driving a fundamental shift in wireless communication systems. Current estimates project over 38 billion IoT connections by 2029, representing a more than twofold increase compared to the 15 million devices in 2023 [1]. One example of such systems is machine-type communications (MTC) (also known as machine-to-machine (M2M) communications), which, as a key aspect of the IoT paradigm, enables machines, sensors, and devices to exchange information without direct human intervention. MTC plays a crucial role in applications ranging from industrial automation and smart cities to healthcare monitoring and environmental sensing [2]. Specifically, massive MTC is anticipated to accommodate a density of up to one million devices (per square kilometer), which typically operate with low computational and storage capabilities, low duty cycles, and small payloads [3]. This massive scale presents unprecedented challenges for communication systems as traditional wireless access paradigms have been designed for a relatively small number of high-data-rate users rather than an enormous number of devices with sporadic low-rate transmissions.
Multiple access (MA) techniques that enable different users to share a common communication medium, have been fundamental to wireless communication systems [4]. MA technologies can be broadly categorized into two categories: centralized access and random access (RA). In centralized access (also known as coordinated access), a central authority controls and manages access to the shared communication channel by allocating distinct resources to each user for data transmission. This approach has been the predominant MA mode in the fifth generation (5G) and previous cellular technologies. However, centralized coordination becomes impractical for massive IoT scenarios due to prohibitive signaling overhead and coordination complexity. In contrast, in RA (also known as uncoordinated access), users transmit their messages without requiring explicit coordination with a central authority. This approach reduces signaling overhead and, in some cases, latency. These advantages make RA solutions particularly suitable for next-generation (6G and beyond) communication networks, especially when serving a large number of users [5, 6].
@@
Traditional RA solutions, such as ALOHA [7] and its variants, suffer from an excessive number of collisions as the number of users increases. More sophisticated grant-free RA schemes [8] employ unique signatures or preambles for each potential user to facilitate collision reduction at the physical layer. Grant-free RA solutions can accommodate more users compared to ALOHA, as they can accomplish user scheduling through a limited amount of feedback [9], or by exploiting device-specific patterns such as codebook structures and spreading sequences [10]. On the other hand, with the number of devices within a certain coverage area reaching millions [15], it is not feasible to assign a distinct pilot/preamble to each device, as this would result in prohibitive receiver complexity. Therefore, existing grant-free RA solutions cannot address the multiple access problem in emerging massive access networks.
@@
To mitigate the challenge of identifying and decoding messages from such a large number of potential devices, Polyanskiy introduced the unsourced random access (URA) paradigm in [11]. URA relies on the observation that, despite their increasing numbers, MTC and IoT devices are activated only sporadically; that is, at any time instant, only a handful of devices are active simultaneously. URA fundamentally reimagines the massive access problem by decoupling user identification from data transmission. Instead of assigning unique preambles or codebooks, users employ a common codebook, and the receiver focuses solely on recovering the list of transmitted messages. Since a common codebook is employed, the user identity is eliminated and the system can operate irrespective of the total number of devices, where only the number of active devices becomes important. This paradigm shift drastically reduces system complexity, and enables practical implementations for truly massive numbers of devices. Following Polyanskiy’s pioneering work, several low-complexity coding solutions have been developed in the subsequent literature, where successive interference cancellation (SIC) techniques are employed to resolve collisions up to a certain level, depending on the user’s transmit powers. Many techniques, such as random spreading [12], multistage orthogonal pilots [13], and time frame slotting [14] have been employed to mitigate collisions and interference. While a single-antenna receiver in a Gaussian MA channel (GMAC) setup is considered in the earlier works, the literature has shifted towards fading channels with multiple-antenna receivers, especially massive multiple-input multiple-output (MIMO), to support more practical channel models and an even greater number of users. A block diagram of different MA techniques is depicted in Fig. 1 alongside the three main setups in URA research, namely, GMAC, and single-antenna and MIMO fading channels.
@@
I-A Objectives and Contributions
Over the years since Polyanskiy’s initial work, a large number of papers have appeared studying URA from both theoretical and practical perspectives. Although prior efforts have been made to review existing works on URA, our survey fills an important gap by providing a systematic classification of the literature and covering more recent works on practical channel models and device characteristics. The first survey on massive access was [15], in which the authors review many-user MA, where each user is assigned an individual codebook (i.e., the coordinated approach), as well as URA, which they call uncoordinated massive access. However, since URA was in its early stages at that time, the survey in [15] covers mainly the earlier works over GMAC. In [16], the focus is again on URA over GMAC, and the authors review the existing design paradigms and coding schemes for URA over GMAC, while URA over fading channels and in MIMO systems are presented as open problems. In a less URA-oriented survey [17], an overview of compressive sensing-based grant-free massive access is provided, which also includes some compressed sensing-based URA schemes. In a recent survey [18], through an overview of the existing URA schemes studying mostly GMAC, the authors present URA as a suitable candidate for next-generation communication systems and highlight that it can support the 3rd Generation Partnership Project (3GPP) standardization efforts.
Figure 2: Structure of the survey.
While the existing surveys focus on the GMAC setup, we also cover the URA solutions that take into account channel fading, which comprise the majority of recent works in the URA literature. More specifically, we conduct a comprehensive review of the existing URA solutions in the literature across three main setups, namely, URA over GMAC, and single-antenna and MIMO fading channels, accompanied by a discussion on the main challenges, possible solutions, and promising research directions. We study works that propose coding schemes, as well as those that provide fundamental information-theoretical analysis for URA. To the best of our knowledge, this is the first comprehensive survey on URA, reviewing this crucial emerging paradigm for 6G and beyond communication networks in all aspects. A summary of the existing surveys and our contribution can be found in Table I.
In particular, the paper contains the following key contributions:
  • Comprehensive overview: Delivering an extensive survey of URA solutions, examining both theoretical and practical perspectives.
  • Classification: Systematically categorizing URA solutions according to the fundamental channel models: GMAC, single-antenna fading channels, and MIMO fading channels.
  • Performance and feasibility analysis: Evaluating and comparing state-of-the-art URA solutions in terms of energy efficiency, computational complexity, and practical viability.
  • Identifying challenges: Addressing essential challenges in URA in terms of interference, collisions, channel estimation, and computational complexity.
  • Guiding future research: Proposing promising future research directions and potential solutions to overcome existing limitations in URA systems.
Throughout the survey, we adopt a unifying narrative: to demonstrate how URA techniques progress from theoretical achievability results on the GMAC to increasingly practical frameworks under fading and MIMO channels. This progression underscores the significance of URA as an enabler of massive connectivity in emerging 6G and IoT systems.
@@ I-A Objectives and Contributions
GB-CR graph-based channel reconstruction and collision resolution
@@
IDMA interleave-division multiple access
@@
MS-MRA multi-stage setup with multiple receive antennas
@@
MSUG multi-stage setup with user grouping
@@
I-B Structure of the Survey
The remainder of the survey is organized as follows. In Section II, a general discussion of MA techniques is provided. We begin with coordinated MA techniques and progress to random access. In Section III, we focus on URA as a new paradigm for uncoordinated MA communications with a massive number of potential devices, enumerate its advantages, and hint at some solutions developed in the literature to address the related challenges. Sections IV and V are dedicated to in-depth overviews of the existing URA solutions, with Section IV focusing on URA over GMAC and Section V on URA over fading channels. Regarding communication models over practical wireless channels, we first investigate the works that take into account channel fading with a single antenna receiver and describe the general approaches; then, we discuss those with a MIMO assumption. Finally, we conclude the survey in Section VI with remarks on URA solutions and provide future research directions. The structure and organization of the paper are illustrated in Fig. 2. Unlike prior surveys that focus primarily on GMAC, our structure integrates both single-antenna and MIMO fading URA solutions, thereby offering a comprehensive view that highlights practical deployment challenges. While our survey spans the full URA literature, the unifying application context is 6G and IoT, where sporadic traffic and massive device density make URA particularly relevant.
@@ II Background: Multiple Access
Multiple access refers to the scenario in which two or more users simultaneously share the system resources. In this section, we provide a concise overview of the main MA approaches and outline their pros and cons to better understand the position of URA solutions in next-generation communication systems.
@@ II-A Coordinated Access
One way to serve multiple users is to assign dedicated resources (e.g., time, frequency, code, or their combinations) to them by a central coordination unit through a prior handshaking procedure. The conventional techniques in this framework include time-division MA (TDMA), frequency-division MA (FDMA), orthogonal frequency-division MA (OFDMA), code-division MA (CDMA), and space-division MA (SDMA). In TDMA and FDMA, non-overlapping time slots and frequency sub-channels are assigned to the different users, respectively, and each user utilizes its allocated time frame or frequency band for transmission [4]. On the other hand, in CDMA, all users can simultaneously use all time-frequency resources through different (nearly) orthogonal code sequences, coupled with a low-complexity decorrelation-based receiver. OFDMA provides frequency division multiplexing by assigning different sets of subchannels to the individual users in an orthogonal frequency division multiplexing (OFDM) system, while SDMA relies on the principle of creating (nearly) orthogonal spatial channels for different users to minimize the inter-user interference by employing beamforming and spatial multiplexing [19].
All the approaches described in the previous paragraph fall under the category of OMA, where the number of users is strictly limited due to orthogonal resource assignment. To increase the number of supported users, non-orthogonal multiple access (NOMA) techniques have also been investigated in the literature, where multiple users can utilize the same resources to transmit their data, thereby allowing the system to support a greater number of users. However, this imposes inter-user interference on the system, and more sophisticated interference-canceling receivers are needed with an increased complexity compared to OMA.
NOMA schemes are divided into two main categories: power-domain NOMA and code-domain NOMA. In power-domain NOMA [20], the users employ different power levels to transmit their data using the same resources, and the receiver exploits the received power differences for detection through SIC. On the other hand, code-domain NOMA is similar to classical CDMA in that users employ user-specific code sequences to transmit their data. The code sequences are sparse or non-orthogonal sequences with low cross-correlations [21]. Here, the receiver can employ message-passing algorithms that utilize the code structure to differentiate between users. Some examples of code-domain NOMA schemes in the literature are low-density spreading code division MA [22], low-density spreading aided orthogonal frequency division multiplexing [23], sparse code MA (SCMA) [24], and multiuser shared access [25].
@@ II-B Random Access
All of the schemes described in the previous subsection are coordinated. That is, they require scheduling grants and coordination with the base station (BS) before transmission, and it is assumed that the number of users and their activity patterns are available at the BS. However, when the number of users becomes large, they become infeasible due to the excessive delays and signaling overhead. For example, the typical number of users per signal dimension in the time-frequency domain (per degree of freedom) is between 1.5 and 3 according to 3GPP studies, as the system performance significantly degrades beyond this threshold [15]. The common solution addressing these drawbacks is random access, where the users perform their transmission without any coordination by the BS whenever they have data. ALOHA and grant-free random access are the two common RA approaches in the literature.
@@ II-B1 ALOHA & Variants
An early RA approach is the ALOHA protocol [7] (also known as pure ALOHA) proposed in the 1970s, where a device sends the entire packet at a randomly selected time. If no acknowledgment is received within a certain period of time, it is assumed that a collision has occurred with a packet sent by another device, and the packet is retransmitted after an additional random waiting time to avoid repeated collisions. This process is repeated until a successful transmission and its subsequent confirmation occur or until the process is terminated by the user console. Slotted ALOHA (SA) is a variant of ALOHA, where packets are transmitted in designated time slots, thereby doubling the maximum throughput compared to pure ALOHA [26]. However, collision is still a significant challenge in SA, as several active users may share the same slot to send their data packets, leading to packets colliding within that slot and the discarding of data. The transmission structures in pure ALOHA and slotted ALOHA are illustrated in Fig. 3.
Figure 3: An illustration of pure and slotted ALOHA. Red boxes show the collided packets, and green boxes are the successfully transmitted ones.
With the motivation to increase the throughput, several variants of SA have been proposed in recent literature, where the main idea is to transmit multiple copies of a packet in different time slots. If the packet is successfully recovered in one slot, the effect of its replicas in the other slots is removed. Based on this idea, two copies of the same packet are transmitted in contention resolution diversity slotted ALOHA (CRDSA) [27], and the packet repetitions are picked from a probability mass function optimized for higher throughput in irregular repetition slotted ALOHA (IRSA), proposed in [28]. An extension of IRSA, exploiting forward error correction, is called coded slotted ALOHA (CSA) [29]. This approach achieves a throughput of almost one when sufficiently low coding rates are employed. IRSA and CSA are further studied in [30, 31, 32, 33, 34]. Moreover, the resolution of partial collisions in random access setups is discussed in [35].